OpenStudy (anonymous):

Could someone help me find the area of this shape? I'm really bad with doing the distance formula.

OpenStudy (igreen):

What shape? @RivS

OpenStudy (anonymous):

Figure ABCDEF is shown. A is located at 5, 6. B is located at 11, 3. C is located at 11, negative 3. D is located at 4, negative 3. E is located at 5, 3. F is located at 1, 3.

OpenStudy (anonymous):

ok where is the shape

OpenStudy (anonymous):

I dont know how to post the shape it wont let me?

OpenStudy (igreen):

Area or perimeter?

OpenStudy (anonymous):

Area

OpenStudy (igreen):

Let's try the triangle, count the units from the bottom to the tip of the triangle

OpenStudy (anonymous):

Its at an angle. How do you count it like that?

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

From the bottom to the top its 3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, now count the base.

OpenStudy (anonymous):

10

OpenStudy (anonymous):

So 1/2bh

OpenStudy (igreen):

Yep

OpenStudy (igreen):

1/2 * 10 * 3

OpenStudy (anonymous):

That would be 1/2 10x3 = 30 = 15

OpenStudy (anonymous):

So the area of the triangle is 15?

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So how do i do the Trapezoid?

OpenStudy (anonymous):

At least I think it's a trapezoid..

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

How would I do it?

OpenStudy (anonymous):

Do I just find the distance between two points?

OpenStudy (anonymous):

I mean two lines

OpenStudy (igreen):

\(A = \dfrac{1}{2}(b_1 + b_2) \times h\)

OpenStudy (igreen):

Nope, we just count units, there's no need to use the distance formula.

OpenStudy (anonymous):

So which one is base 1?

OpenStudy (anonymous):

Ok

OpenStudy (igreen):

The one on the top, but it doesn't really matter

OpenStudy (anonymous):

A = 1/2 (7 + 6) x 6

OpenStudy (anonymous):

A = 1/2 (13) x 6

OpenStudy (anonymous):

A = 1/2 78

OpenStudy (anonymous):

So the area is 39?

OpenStudy (anonymous):

39 + 15 = 54

OpenStudy (anonymous):

@iGreen Thanks for the help

OpenStudy (igreen):

Hold on

OpenStudy (igreen):

Yes, excellent work! \(\Huge\color{lime} \checkmark\)